OpenStudy (anonymous):

if m<4 = (3x)o and m<8 = (x+40)o what is the measure of <4?

OpenStudy (anonymous):

I think more information is needed.

OpenStudy (anonymous):

thats all thats there expect to pictures

OpenStudy (anonymous):

60

OpenStudy (anonymous):

<4 = <8 3x = x + 40 2x = 40 x = 20 m<4 = 3(20) = (60) o

OpenStudy (anonymous):

So these angles are equal? How do you make that assumption? There is not enough info given to solve for x.